The manager of an insurance company wants to know the correlation number of customer services representatives and the number of returning insured each year. The results are as follows:

# of reps # of returning insureds (100s)

1

2

3

5

8

11

11

17

20

25

5

10

15

17

18

20

20

26

27

30

Find the correlation between customer service representatives and returning insured. ??
Answer #1

Solution:

Given :

The manager of an insurance company wants to know the correlation number of customer services representatives and the number of returning insured each year. The results are as follows:

# of reps # of returning insureds (100s)

1

2

3

5

8

11

11

17

20

25

5

10

15

17

18

20

20

26

27

30

Correlation coefficient formula is :

r=\frac{SS_{xy}}{\sqrt{SS_{xx}\times SS_{yy}}}

where

SS_{xy}=\sum xy \: \: -\: \: \left ( \sum x\times \sum y \: \: /\: \: n\: \: \right )

SS_{xx}=\sum x^{2} \: \: -\: \: \left ( \sum x\times \sum x \: \: /\: \: n\: \: \right )

SS_{yy}=\sum y^{2} \: \: -\: \: \left ( \sum y\times \sum y \: \: /\: \: n\: \: \right )

Thus we need to make following table:

x : # of reps y : # of returning
insureds (100s)
x^2 y^2 xy
1 5 1 25 5
2 10 4 100 20
3 15 9 225 45
5 17 25 289 85
8 18 64 324 144
11 20 121 400 220
11 20 121 400 220
17 26 289 676 442
20 27 400 729 540
25 30 625 900 750
\sum x = 103 \sum y = 188 \sum x^{2} = 1659 \sum y^{2} = 4068 \sum xy= 2471

Thus

SS_{xy}=\sum xy \: \: -\: \: \left ( \sum x\times \sum y \: \: /\: \: n\: \: \right ).

SS_{xy}=2471 \: \: -\: \: \left ( 103\times 188 \: \: /\: \: 10\: \: \right )

SS_{xy}=2471 \: \: -\: \: 1936.4

SS_{xy}=534.6

SS_{xx}=\sum x^{2} \: \: -\: \: \left ( \sum x\times \sum x \: \: /\: \: n\: \: \right )

SS_{xx}=1659 \: \: -\: \: \left ( 103 \times 103 \: \: /\: \: 10\: \: \right )

SS_{xx}=1659 \: \: -\: \: 1060.9

SS_{xx}=598.1

SS_{yy}=\sum y^{2} \: \: -\: \: \left ( \sum y\times \sum y \: \: /\: \: n\: \: \right )

SS_{yy}=4068 \: \: -\: \: \left ( 188 \times 188 \: \: /\: \: 10\: \: \right )

SS_{yy}=4068 \: \: -\: \: 3534.4

SS_{yy}=533.6

Thus correlation coefficient is :
r=\frac{SS_{xy}}{\sqrt{SS_{xx}\times SS_{yy}}}

r=\frac{534.6 }{\sqrt{598.1 \times 533.6 }}

r=\frac{534.6 }{\sqrt{319146.16 }}

r=\frac{534.6 }{564.9302258 }

r=0.9463

Thus the correlation between customer service representatives and returning insured is r = 0.9463 .